find the bases of a trapezoid if the height is 10,the area is 60,and one base is twise the other what is the answer?

Answer by mananth(16946)

one parallel side =x
other parallel side=2x
height =10
area = 60
Area of trapezoid = 1/2(sum of parallel sides ) *height
60=1/2 * (x+2x)*10
60=1/2*3x*10
60=15x
x=4
one side = 4
other side = 8